For the sentence y =x -2
X 2 5 9
Y ? ? ?
I don't understand ???

Answer by josgarithmetic(39630)

--------------------------------
For the sentence y =x -2 

X 2 5 9 
Y ? ? ? 

I don't understand ??? 

-------------------------------

You are expected to make a table of values for x and y, for the EQUATION y=x-2. THAT is what it means.
Equation:  y=x-2

x |  2   5   9
---------------
y |